OpenStudy (anonymous):

Find the volume of a cube measuring 1.5 cm on each side.

OpenStudy (anonymous):

cube it hence the name \[(1.5)^3\]

OpenStudy (anonymous):

^ that made me laugh

OpenStudy (anonymous):

if it was a square, you would square it if it was a circle, you would circle it

OpenStudy (anonymous):

um… so would it be 4.5?

OpenStudy (anonymous):

Volume of a cube is side cubed

OpenStudy (anonymous):

no don't multiply it by 3 cube it

OpenStudy (anonymous):

\[(1.5)^3 = (1.5)(1.5)(1.5)\]

OpenStudy (anonymous):

\[(1.5)^3=1.5\times 1.5\times 12.5\]

OpenStudy (anonymous):

oh ya duh!!! sorry!

OpenStudy (anonymous):

3.375?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thanks!